OpenStudy (anonymous):

s4 – 16 A. (s - 2)2(s + 2)2 B. (s - 2)(s + 2) C. (s - i)(s + i)(s - 2)(s + 2) D. (s - 2i)(s + 2i)(s - 2)(s + 2)

OpenStudy (anonymous):

I think the answer is B

OpenStudy (anonymous):

Again, no. Do you recognize that this is a difference in squares?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok. Do you see how s^2 * s^2 = s^4? and that 4 * 4 = 16?

OpenStudy (anonymous):

yes so that means the answer has to be A

OpenStudy (anonymous):

Nope... \[s^{4} - 16 = (s^{2} - 4)(s^{2} + 4)\]Now, do you realize that s^2 - 4 is also a difference in squares?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Alright. The form \[a^{2} - b^{2} = (a + b)(a - b)\]Therefore, \[s^{2} - 4 = (s + 2)(s - 2) The same for s^2 + 4. It turns imaginary. \[s^{2} + 4 = (s - 2i)(s + 2i)

OpenStudy (anonymous):

so the answer is C

OpenStudy (anonymous):

You've guessed every answer except the correct one...

OpenStudy (anonymous):

ok thanks for the help I get it now.

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

What is the degree of each monomial? 7m6n5 A. 5 B. 11 C. 6 D. 7

OpenStudy (anonymous):

how do you firgure out the degrees?